ou want to buy a car, and a local bank will lend you $10,000. The loan would be fully amortized over 5 years (60 months), and the nominal interest rate would be 15%, with interest paid monthly. What is the monthly loan payment? Do not round intermediate calculations. Round your answer to the nearest cent.

$  

What is the loan's EFF%? Do not round intermediate calculations. Round your answer to two decimal places.

%
Answer #1
monthly loan repayment= P*R*(1+R)^N]/[(1+R)^N-1
where, P= Principal= $10,000, R= monthly rate of interest=1.25% (15%/12), N= number of payments= 60 (5*12 months)

monthly loan payment=$10,000*0.0125*(1+0.0125)^60/[(1+0.0125)^60]-1

Monthly repayment is $237.8993 or $238

Monthly interest payment= R/N*P = (0.15/12)* balance principal outstanding at the beginning of the year

where, P= Principal at the start of the year, R= rate of interest, N= number of payments i.e 12 times in a year

Effective annual interest rate = (1 + (nominal rate / number of compounding periods)) ^ (number of compounding periods) - 1
where, nominal rate= 15%, number of compounding period= 12 i.e 12 months
effective interest rate= (1+(0.15/12))^12-1

effective interest rate= 16.0755%

Below have attached the repayment schedule for your reference.

Month Starting balance Repayment Interest paid Principal paid New balance
1 $10,000 237.899 $125 $113 $9,887
2 $9,887 237.899 $124 $114 $9,773
3 $9,773 237.899 $122 $116 $9,657
4 $9,657 237.899 $121 $117 $9,540
5 $9,540 237.899 $119 $119 $9,421
6 $9,421 237.899 $118 $120 $9,301
7 $9,301 237.899 $116 $122 $9,179
8 $9,179 237.899 $115 $123 $9,056
9 $9,056 237.899 $113 $125 $8,932
10 $8,932 237.899 $112 $126 $8,805
11 $8,805 237.899 $110 $128 $8,678
12 $8,678 237.899 $108 $129 $8,548
13 $8,548 237.899 $107 $131 $8,417
14 $8,417 237.899 $105 $133 $8,284
15 $8,284 237.899 $104 $134 $8,150
16 $8,150 237.899 $102 $136 $8,014
17 $8,014 237.899 $100 $138 $7,876
18 $7,876 237.899 $98 $139 $7,737
19 $7,737 237.899 $97 $141 $7,596
20 $7,596 237.899 $95 $143 $7,453
21 $7,453 237.899 $93 $145 $7,308
22 $7,308 237.899 $91 $147 $7,161
23 $7,161 237.899 $90 $148 $7,013
24 $7,013 237.899 $88 $150 $6,863
25 $6,863 237.899 $86 $152 $6,711
26 $6,711 237.899 $84 $154 $6,557
27 $6,557 237.899 $82 $156 $6,401
28 $6,401 237.899 $80 $158 $6,243
29 $6,243 237.899 $78 $160 $6,083
30 $6,083 237.899 $76 $162 $5,921
31 $5,921 237.899 $74 $164 $5,757
32 $5,757 237.899 $72 $166 $5,591
33 $5,591 237.899 $70 $168 $5,423
34 $5,423 237.899 $68 $170 $5,253
35 $5,253 237.899 $66 $172 $5,081
36 $5,081 237.899 $64 $174 $4,906
37 $4,906 237.899 $61 $177 $4,730
38 $4,730 237.899 $59 $179 $4,551
39 $4,551 237.899 $57 $181 $4,370
40 $4,370 237.899 $55 $183 $4,187
41 $4,187 237.899 $52 $186 $4,001
42 $4,001 237.899 $50 $188 $3,813
43 $3,813 237.899 $48 $190 $3,623
44 $3,623 237.899 $45 $193 $3,431
45 $3,431 237.899 $43 $195 $3,236
46 $3,236 237.899 $40 $197 $3,038
47 $3,038 237.899 $38 $200 $2,838
48 $2,838 237.899 $35 $202 $2,636
49 $2,636 237.899 $33 $205 $2,431
50 $2,431 237.899 $30 $208 $2,223
51 $2,223 237.899 $28 $210 $2,013
52 $2,013 237.899 $25 $213 $1,800
53 $1,800 237.899 $23 $215 $1,585
54 $1,585 237.899 $20 $218 $1,367
55 $1,367 237.899 $17 $221 $1,146
56 $1,146 237.899 $14 $224 $923
57 $923 237.899 $12 $226 $696
58 $696 237.899 $9 $229 $467
59 $467 237.899 $6 $232 $235
60 $235 237.899 $3 $235 ($0)